OpenStudy (chycora):

Determine the equation of the line that passes through the points (2, -5) and (7, 0). Use the form y = mx + b. A. y = x + 7 B. y = 5x + 5 C. y = 2x - 5 D. y = x - 7

OpenStudy (mrnood):

do oyu understand the form of the equation? m is the slope of the line and b is the point where it crosses the y axis are oyu ok with this for a start?

OpenStudy (chycora):

yes

OpenStudy (mrnood):

do you know a formula for 'slope'? your points are x1=2 y1= -5 x2=7 y2= 0

OpenStudy (chycora):

so would the answer be c

OpenStudy (mrnood):

how did oyu get that? the formula for slope is \[slope = \frac{ y _{2} - y _{1} }{ x _{2} -x _{1} }\] Use the values I gave you above and that equation to work out what the slope is

OpenStudy (chycora):

i'm working it out

OpenStudy (chycora):

the answer is b right

OpenStudy (mrnood):

work out the slope

OpenStudy (chycora):

i did

OpenStudy (mrnood):

What value did you get for slope?

OpenStudy (chycora):

\[slope =\frac{-5 }{ 5 }\]

OpenStudy (chycora):

that's wat i gt

OpenStudy (mrnood):

that is not quite correct what is 0 - (-5)?

OpenStudy (chycora):

5

OpenStudy (mrnood):

so what is the slope?

OpenStudy (chycora):

y = 5x + 5

OpenStudy (mrnood):

Please just answer the questions - they are guiding you to work out the answer properly you said the slope was -5/5 but that is not correct you said 0-(-5) = 5 which IS correct so just answer 'what is the slope'?

OpenStudy (mrnood):

B is not correct